OpenStudy (anonymous):

Given the function f(x) = 0.5(3)^x, what is the value of f−1(7)?

OpenStudy (isaiah.feynman):

First find the inverse.

OpenStudy (jdoe0001):

\(\bf f(x) = y= 0.5(3)^x\qquad inverse\implies x= 0.5(3)^y\) solve for "y"

OpenStudy (anonymous):

i have no idea how to do this i need some explaining :(

OpenStudy (jdoe0001):

notice, all you do is swap the variables about, to get the inverse

OpenStudy (anonymous):

oh okay i understand

OpenStudy (anonymous):

what do i need to do now? @jdoe0001

OpenStudy (anonymous):

so now i have x= 0.5(3)^y how do i find the value of f^−1(7)?

OpenStudy (jdoe0001):

hmm... have you cover logarithms yet?

OpenStudy (anonymous):

no, just exponential functions..

OpenStudy (jdoe0001):

hmm... well.... then maybe that's all, you leave it as exponential because to solve for "y", it'd end up as a logarithmic function

OpenStudy (anonymous):

okay, thats it?

OpenStudy (jdoe0001):

I'd think so, yes :)

OpenStudy (anonymous):

that not one of the options :/

OpenStudy (anonymous):

Given the function f(x) = 0.5(3)x, what is the value of f−1(7)? 0.565 1.140 1.771 2.402

OpenStudy (jdoe0001):

hmmm so what are your choices?

OpenStudy (jdoe0001):

ohhh... wait.. your options will be just a value... however... to get that value, you'd need the logarithmic function, which you haven't covered

OpenStudy (anonymous):

okay thanks :) ill try and figure it out